OpenStudy (anonymous):

what is the surface area of a circle that has a diameter or 14 inches?? FAST PLEASE

OpenStudy (opcode):

Formula: \(area = π · r2\)

OpenStudy (opcode):

Formula: \(area = π \times r^2\) Radius = r Radius is half of the diameter. But since we are squaring it. We don't need to divide the diameter in half. Plug in the numbers: \(area = 3.14 \times 14\) Answer: 43.96